OpenStudy (anonymous):

Use the Substitution Method to solve the following system of equations. 2x - 9y = 1 x - 4y = 1

OpenStudy (anonymous):

HELPPPPPPPPPPPPPPPP

OpenStudy (anonymous):

start with \[x=1+4y\] the first equation becomes \[2(1+4y) - 9y = 1 \]

OpenStudy (anonymous):

x=1+4y 2(1+4y)-9y=1 2+8y-9y=1 y=1

OpenStudy (anonymous):

thank you all